A girl pulls her school box with a force of 78 Newton inclined at 27 degrees to the ground calculate the part of the force moving the box along the ground
1
Expert's answer
2019-11-05T10:39:00-0500

The horizontal component of the force


"F_x=F\cos 27\degree=78\:\rm N\times \cos 27\degree=69.5\: N"